Upton Umbrellas has a cost of equity of 12.1 percent, the YTM on the company's bonds is 6.2 percent, and the tax rate is 39 percent. The company's bonds sell for 103.7 percent of par. The debt has a book value of $423,000 and total assets have a book value of $957,000. If the market-to-book ratio is 2.89 times, what is the company's WACC?

  • 9.96%

  • 5.62%

  • 8.49%

  • 10.26%

  • 8.35%
Answer #1

Book Value of Equity = Total Assets - Book Value of Debt

Book Value of Equity = 957,000 - 423,000 = $534,000

Market Value of Equity = 534,000(2.89) = $1,543,260

Market Value of Debt = 423,000(1.037)

Market Value of Debt = $438,651

WACC = [438,651(0.062)(1 - 0.39) + 1,543,260(0.121)]/(438,651 + 1,543,260)

WACC = 0.1026

WACC = 10.26%